using this data 3,13,15,16,16,20,23 N=7 what score would a z of -.25 be and what score would a z of 1.5 be. please show all work.
Answer by stanbon(75887)

using this data 3,13,15,16,16,20,23 N=7 what score would a z of -.25 be and what score would a z of 1.5 be. please show all work.
--------------
Find the mean of the data::
(3+13+15+16+16+20+23)/7 = 106/7 = 15.14
----
std = 5.84
-----
score = z*s+u
Ans: score = -0.25*5.84+15.14 = 13.6
---------------
Cheers,
Stan H.
--------------